# The Repeat Offenders: Characterizing and Predicting Extremely Bug-Prone Source Methods

### v1

@@ 3.2. Labeling Bug-fix Commits
The bug-proneness of a method is historically estimated by extracting bug-related keywords from commit messages. A commonly used set of such keywords (and their variations) includes: error, bug, fix, issue, mistake, incorrect, fault, defect, and flaw (ray_naturalness_2016; mockus_identifying_2000). However, this keyword set often produces false positives, primarily due to the inclusion of the keyword issue (chowdhury_method-level_2024). To reduce these false positives, Chowdhury et al. (chowdhury_method-level_2024) proposed removing the keyword issue and focusing on commit messages that contain both bug-related and fix-related keywords. Although this method significantly improves the accuracy of identifying bug-fix commits, it does not eliminate noise caused by tangled commits—commits that include changes to both bug-related and unrelated methods—leading to incorrectly labeled methods.
While handling tangled code changes remains an active research area (herbold2022fine; opu2025llm; chowdhury_method-level_2024), we conduct our analysis using three distinct datasets to mitigate the impact of tangled commits. Our underlying hypothesis is that if similar results are observed across all three datasets, the findings are more robust and less likely to be impacted by tangled changes.
@@
    HighRecall Dataset. Here, we used all the above-mentioned keywords and their variations (include issue) and did not consider the problem of tangled changes. While this approach obtains high recall since any change remotely related to a bug-fix is labeled as a bug-fix, it results in low precision as many changes not related to a bug-fix are labeled as such.
@@
    HighPrecision Dataset. In this dataset we used the set of keywords introduced by Chowdhury et al. (chowdhury_method-level_2024), which aims at increasing the precision of bug-fix labeling. To remove the effects of tangled changes, we only labeled a change as a bug-fix if at most 1 method was changed in that commit—if there is only one modified method in a bug-fix commit, that method was certainly a bug-prone method (chowdhury_method-level_2024). While this approach achieves a high level of precision, it ultimately reduces recall, as many bug fixes may be ignored.
@@ 3.3. Age Normalization
The age of the methods in our dataset varies significantly. This means that comparing newer methods to older methods may introduce bias, as older methods have more time to undergo changes and hence, undergo bug-fix related changes. This assumption is confirmed in figure 1(a), where we can see that for many projects, there is a positive correlation when comparing age against both revisions and bugs. Specifically, for approximately 60% of projects, the correlation between a method’s age and its number of bugs is $\geq 0.2$—while weak, such correlations still should not be ignored to reduce confounding in our analysis.
Due to this correlation, we removed all methods that are under 5 years old from our dataset. However, we still cannot directly compare methods that are 5 years old with those that are more than 5 years old for the aforementioned reasons. Thus, we also remove all changes made after the first 5 years from the remaining methods. This, unfortunately, resulted in us losing 626,529 methods and all methods from the following 6 projects from our dataset: ’junit5’, ’SmartTube’, ’RxJava’, ’Essentials’, ’titan’, ’xpipe’. Since we are interested in the Extremely Bug-Prone methods, we care more about retaining change history—allowing methods to exhibit their true bug-proneness. Thus, the reasoning behind a 5-year threshold can be summarized in Figure 1(b). Here we see that 5 years allows us to retain ~90% of changes and ~85% of bugs while still saving ~50% of the methods in our dataset. If we decrease the threshold, we will lose more bug-proneness information in our dataset.
The end result of this age-normalization is a dataset of 691,610 methods from 92 projects, where each method has exactly 5 years of change history. All the projects in our dataset, after age normalization, as well as the number of ExtremelyBuggy methods in each project for the Balanced dataset can be viewed in our public repository  [^8] https://github.com/SQMLab/ExtremelyBuggyPublicData/tree/main/AdditionalFigures/RepositoryList.md.
(a) Correlation between age and both revisions and bug-proneness for all 98 projects
(b) Percent of total methods, revisions and bugs retained at different age-normalization thresholds.
Figure 1. Both figures use the HighRecall Dataset to measure bugs. To enhance readability, we only use 11 markers when presenting each distribution, however, all data points are represented.:
@@ 3.4. Labeling Bug Proneness
For all three datasets, we assign each method to one of three categories based on the number of bug-fix commits in which it was involved:
@@
3.5. Code Metrics
Code metrics have traditionally served as indicators of quality and maintainability of code (gil_correlation_2017; pascarella_performance_2020; chowdhury_method-level_2024). Encouraged by previous research in method-level code metrics (mo2022exploratory; pascarella_performance_2020; chowdhury_evidence_2025), we used 14 code metrics to determine if the code quality indicators of the ExtremelyBuggy methods are different than others. The selected metrics are described as follows.
Size. Size is one of the simplest metrics, yet it is considered one of the most important with regard to software quality and maintenance (gil_correlation_2017; el2001confounding; chowdhury2022empirical). We used the number of source lines of code without comments and blank lines as our size measurement (referred to as SLOCStandard) (chowdhury_method-level_2024; landman2014empirical; ralph2018construct; chowdhury_evidence_2025).
Readability. The ability to read and understand existing code is an important part of software maintenance (scalabrino2016improving). Code readability can be broadly defined by how easy it is for a developer to understand the structure and flow of source code (buse2009learning; posnett2011simpler; borstler_role_2016; johnson2019empirical). Developers can more easily understand and therefore modify readable code with reduced risk of issues. We used two distinct readability metrics in this paper: Buse et al. (buse2009learning) (Readability) and Posnett et al. (posnett2011simpler) (SimpleReadability). Both provide a value for each method in the range $[0,1]$.
Complexity and Testability. We used the McCabe metric (mccabe1976complexity) to analyze the complexity of methods. This metric represents the number of independent paths in a method; the greater the number of paths, the more challenging it becomes to test the method thoroughly. In addition, we used the number of control variables (NVAR), the number of comparisons (NCOMP), and the McClure metric (mcclure1978model), which is calculated as the sum of NVAR and NCOMP. These metrics offer additional insights into method complexity that are not captured by the McCabe metric. Hindle et al. (hindle2008reading) proposed the proxy indentation metric (IndentSTD), which performs similarly to the McCabe metric in capturing code complexity but, unlike McCabe, does not require a language-specific parser. We included IndentSTD in our metric set as well. Finally, we used the MaximumBlockDepth metric, which measures the maximum nested block depth, as deep nesting within a method can further complicate testing.
Maintainability. Maintainability Index (MI) (oman1992metrics) is a composite software quality metric that integrates HalsteadVolume, which quantifies the amount of information a reader must process to comprehend the code’s functionality, along with McCabe complexity and Size. The MI is computed using the following formula:
Other. Halstead Length (Length), which measures the total number of operators and operands, the number of parameters (Parameters) and the number of local variables (LocalVariables), are also included alongside the other metrics. We also intended to include the FanIn metric in our analysis, but were unable to do so due to the complexity involved. For example, consider the Hadoop project, which contains approximately 70,000 methods. Since our goal is to make predictions at method inception, we need to capture code metrics at the time each method was first introduced. Measuring the FanIn metric would require constructing a separate call graph for the whole codebase at each method’s introduction commit. Unfortunately, building such call graphs is extremely time-consuming, as it requires preprocessing the entire codebase. Given that we work with 98 projects—each potentially requiring thousands of unique call graphs corresponding to thousands of distinct commits—this approach was computationally infeasible within the scope of our study.
@@ 3.6. Statistical Tests
After evaluating our data with the Anderson-Darling normality test (razali2011power), we found that most of the distributions in our data (e.g., the code metrics) do not follow a normal distribution. Therefore, we adopted the non-parametric Wilcoxon rank sum test to assess if there is a statistical difference between two given distributions. To understand the size of those difference, we employed the non-parametric Cliff’s d and categorized the effect sizes following Hess et al. (hess2004robust): Negligible (N) (<0.147<0.147), Small (S ($0.147\leq\delta<0.33$), Medium (M) ($0.33\leq\delta<0.474$), and Large (L) ($\delta\geq 0.474$). Both of these statistical tests do not assume normally distributed distributions and have been widely used in software engineering research (he2015empirical; chen2020savior; pecorelli2020testing; chowdhury2019greenscaler; bangash2020time). In addition, we used the non-parametric Kendall’s $\tau$ correlation coefficient instead of Pearson’s r (inozemtseva2014coverage; gil_correlation_2017; chowdhury2019greenscaler). Unless otherwise stated, we use a significance threshold of $p\leq 0.05$ when evaluating the statistical significance of correlation coefficients.
@@ 4.1. RQ1: Proportion of ExtremelyBuggy methods and their impact
Past research has shown that a high proportion of changes come from a small segment of code (grosfeld2007pareto; chowdhury_good_2025). Motivated by this observation, we investigate what proportion of methods are ExtremelyBuggy and how many bugs the ExtremelyBuggy methods account for in a project.
@@
Table 1 shows the number of Buggy and ExtremelyBuggy methods for all three different datasets when we consider all 92 projects aggregated together. Although the high recall dataset contains a large number of ExtremelyBuggy methods (6.63%), these numbers decline significantly when we observe the HighPrecision and Balanced datasets. Specifically, only 0.04% (or 287 of the ~700000) methods are ExtremelyBuggy in the HighPrecision dataset. Another observation is that the ratio of Buggy to ExtremelyBuggy methods is smaller in the HighPrecision and balanced datasets. In the HighRecall dataset, approximately 37% of all methods with at least one bug-fix (Buggy and ExtremelyBuggy) are ExtremelyBuggy whereas in the HighPrecision and Balanced datasets, this percentage drops to approximately 10% and 13% respectively.
By viewing the results in figure 2, we can see the percent of methods that are Buggy and ExtremelyBuggy for individual projects across all datasets. Each plot shows the CDF of the percent of methods that are Buggy and ExtremelyBuggy on a log scale where each data point is an individual project. For example, in the HighPrecision dataset, in ~80% of projects, $\leq 1\%$ of methods are Buggy. One interesting finding is that for the HighPrecision dataset, ~35% of projects have no ExtremelyBuggy methods. This will make it very hard to do a project-wise analysis on the ExtremelyBuggy methods using this dataset since the results will be skewed by the large number of projects with zero ExtremelyBuggy methods.
@@
Figure 2. CDFs of the percent of methods in that are Buggy and ExtremelyBuggy for each of the 3 datasets for all projects. To enhance readability, we only use 11 markers when presenting each distribution, however, all 92 projects are represented.:
Figure 3. CDFs of the percent of bugs covered by the ExtremelyBuggy methods for each of the three datasets for all 92 projects, excluding projects with no ExtremelyBuggy methods. To enhance readability, we only use 11 markers when presenting each distribution.:
We now investigate what percent of bugs are captured by the ExtremelyBuggy methods alone. We counted the number of bugs in the ExtremelyBuggy methods in each project and divided it by the total number of bugs in the project for each of the 92 projects. To avoid overcounting the number of bugs in each project, we only count a bug once for each unique commit. The results for each dataset are presented in figure 3. For the HighRecall Dataset, we can see that the ExtremelyBuggy methods account for a high proportion of bugs, as for ~80% of projects the ExtremelyBuggy methods accounted for $\geq 60\%$ of bugs and ~50% of projects the ExtremelyBuggy methods covered $\geq 75\%$ of bugs. Not surprisingly, the results are much different for the HighPrecision and Balanced datasets. For these two datasets, we see much less bug coverage of the top percent of methods with bugs. However, this is expected as in the HighPrecision and Balanced datasets, many of the ExtremelyBuggy methods were excluded for maintaining higher precision.
RQ1 Summary: Across all datasets, only a small fraction of methods are ExtremelyBuggy, yet they account for a disproportionately large share of bugs. In the HighRecall dataset, these rare methods account for a large share of bugs, often more than 60% for 80% of the projects. This underscores the importance of early identification and remediation of these methods, as proactive action is crucial for minimizing future maintenance costs.
@@ 4.2. RQ2: Code Quality of ExtremelyBuggy Methods
Figure 4. CDFs showing the distributions of different code metrics at the first commit between NotBuggy, Buggy, and ExtremelyBuggy methods for the Balanced dataset. To enhance readability, we only use 11 markers when presenting each distribution; however, all methods are represented. Note that 34 of the ~700000 methods had a MaintainabilityIndex below zero. To improve readability, we excluded these methods when constructing the MaintainabilityIndex graph.:
Figure 4 compares the three different method classes for six of the selected code metrics using the Balanced dataset; the results are similar for all code metrics across all three datasets. Clearly, ExtremelyBuggy methods are larger, less readable, and more complex than both the Buggy and NotBuggy methods.
To confirm these visual findings, we also performed a statistical analysis to compare ExtremelyBuggy methods with both Buggy and NotBuggy methods. In Table 2 we present the results when comparing the ExtremelyBuggy methods to Buggy and NotBuggy methods grouped together for the Balanced dataset. The comparisons were done on all 92 projects’ distributions aggregated together. We can see that all the P-values are significant ($P\leq 0.05$) and that all the effect sizes are large, except for Parameters, which is medium. This indicates that there is a substantial difference between ExtremelyBuggy methods and all other methods at inception. Additionally, the sign of the effect size is negative (-) for Readability, SimpleReadability, and MaintainabilityIndex indicating that these methods are less readable and maintainable. For other code metrics the sign is positive (++) indicating that the ExtremelyBuggy methods are larger and more complex. The results for the HighRecall dataset are the same; however, the effect sizes for the HighRecall dataset are mostly medium. This is because the HighRecall dataset has a higher number of false positives when identifying the ExtremelyBuggy methods, resulting in more noise and less distinction of the ExtremelyBuggy methods.
Table 2. The results of statistical significance test (Wilcoxon rank sum test) comparing ExtremelyBuggy methods with the combined group of Buggy and NotBuggy methods for the Balanced dataset dataset. The method groups used in the comparison are formed by aggregating methods from all projects.
@@
In addition to the aggregated analysis, we also perform individual project analysis to obtain generalizable observations. Projects that has no ExtremelyBuggy methods were excluded from this analysis. Table 3 presents the results for the ExtremelyBuggy compared to the Buggy and NotBuggy methods grouped together for the Balanced dataset. For example, for 78.21% of projects, the difference in SLOCStandard between ExtremelyBuggy and other methods is statistically significant ($P\leq 0.05$). For these projects, the difference was large (L) for 98.36% of the projects and for the remaining 1.64% it was medium (M). For all code metrics, some projects (usually approximately 20-30%) did not show statistical differences between ExtremelyBuggy and other methods. This is because some projects in the Balanced dataset have a very low number of ExtremelyBuggy methods, making it difficult to discern differences in distributions. However, for the projects that did have a statistical difference, that difference was almost always large. Out of the 14 code metrics, only one, Parameter, exhibited distinct characteristics. While more than 62% of projects consistently demonstrated statistical significance for all other metrics, Parameter differed for only 20% of projects, indicating that the bug-proneness of methods is not significantly influenced by parameter count.
For the HighPrecision dataset, approximately 40% of projects exhibited statistical differences between ExtremelyBuggy and other methods for all metrics, with the effect size remaining almost always large. For the HighRecall dataset, approximately 90-95% of projects demonstrated statistical differences for all metrics; however, the effect size ranged from small (S) to large (L) (typically ~30% in each category). This is expected due to the high number of ExtremelyBuggy methods but lower precision in identifying ExtremelyBuggy methods present in the HighRecall dataset.
Table 3. The results of statistical significance test (Wilcoxon rank sum test) and effect size (Cliff’s delta) comparing ExtremelyBuggy methods with the combined group of Buggy and NotBuggy methods for the Balanced dataset dataset, evaluated per project. Each value denotes the percentage of projects exhibiting the corresponding behavior. For instance, for the totalFanOut metric, 80.77% of projects show a statistically significant difference, and in 98.41% of those cases, the effect size is classified as large (L).
@@
To further understand how ExtremelyBuggy methods compare to Buggy and NotBuggy ones, we also did these comparisons without grouping the Buggy and NotBuggy methods together. We found that the effect size differences between NotBuggy and Buggy methods were generally small to medium and the same was found when comparing the Buggy to ExtremelyBuggy methods. However, the difference between NotBuggy and ExtremelyBuggy methods was mostly large for all code metrics and datasets. This finding confirms the visual findings from figure 4 about the differences between the three bug-proneness types.
RQ2 Summary: At their inception, ExtremelyBuggy methods are significantly larger, less readable, and more complex than Buggy and NotBuggy methods. This is encouraging because it suggests these methods are distinguishable at the very beginning of their lifetime.

### v2

@@ 3.3. Labeling Bug-fix Commits
The bug-proneness of a method is historically estimated by extracting bug-related keywords from commit messages. A commonly used set of such keywords (and their variations) includes: error, bug, fix, issue, mistake, incorrect, fault, defect, and flaw (Ray et al., 2016; Mockus and Votta, 2000). However, this keyword set often produces false positives, primarily due to the inclusion of the keyword issue (Chowdhury et al., 2024). To reduce these false positives, Chowdhury et al. (Chowdhury et al., 2024) proposed removing the keyword issue and focusing on commit messages that contain both bug-related and fix-related keywords. Although this method significantly improves the accuracy of identifying bug-fix commits, it does not eliminate noise caused by tangled commits—commits that include changes to both bug-related and unrelated methods—leading to incorrectly labeled methods.
While handling tangled code changes remains an active research area (Herbold et al., 2022; Opu et al., 2026; Chowdhury et al., 2024), we conduct our analysis using three distinct datasets to mitigate the impact of tangled commits. The underlying hypothesis is that if similar results are observed across all three datasets, the findings are more robust and less likely to be impacted by tangled changes (Chowdhury et al., 2025).
@@
    HighRecall Dataset. Here, we used all the above-mentioned keywords and their variations (include issue) and did not consider the problem of tangled changes. While this approach obtains high recall since any change remotely related to a bug-fix is labeled as a bug-fix, it results in low precision as many changes not related to a bug-fix are labeled as such. Since this dataset is unlikely to miss any bugs, it is well-suited for estimating the proportion of bugs that can be identified by considering only the ExtremelyBuggy methods (RQ1).
@@
    HighPrecision Dataset. In this dataset, we used the set of keywords introduced by Chowdhury et al. (Chowdhury et al., 2024), which aims at increasing the precision of bug-fix labeling. To remove the effects of tangled changes, we only labeled a change as a bug-fix if exactly one method was changed in that commit—if there is only one modified method in a bug-fix commit, that method was certainly a bug-prone method (Chowdhury et al., 2024). While this approach achieves a very high level of precision, it ultimately reduces recall, as many bug fixes may be ignored. Because this dataset is small and precise, it is particularly well suited for the thematic analysis (RQ4) aimed at understanding the properties of ExtremelyBuggy methods.
@@ 3.4. Age Normalization
The age of the methods in our dataset varies significantly. This means that comparing newer methods to older methods may introduce bias, as older methods have more time to undergo changes and hence, undergo bug-fix related changes. This assumption is supported by Figure 2(a), which shows that for many projects there is a positive correlation—measured using Kendall’s $\tau$ correlation coefficient with $p\leq 0.05$—between age and both revisions and bugs. Specifically, for approximately 60% of projects, the correlation between a method’s age and its number of bugs is $\geq 0.2$. While weak, such correlations should still be considered to minimize potential confounding in our analysis.
Due to this correlation, we removed all methods that are under 5 years old from our dataset. However, we still cannot directly compare methods that are 5 years old with those that are more than 5 years old for the aforementioned reasons. Thus, we also remove all changes made after the first 5 years from the remaining methods. This, unfortunately, resulted in us losing 593,409 methods and all methods from the following 6 projects from our dataset: ’junit5’, ’SmartTube’, ’RxJava’, ’Essentials’, ’titan’, ’xpipe’. Since we are mainly interested in ExtremelyBuggy methods, we care more about retaining change history—allowing methods to exhibit their true bug-proneness. Thus, the reasoning behind a 5-year threshold can be summarized in Figure 2(b). Here we see that 5 years allows us to retain ~90% of changes and ~85% of bugs while still saving ~50% of the methods in our dataset. If we decrease the threshold, we lose more bug-proneness information in our dataset, which subsequently reduces the number of ExtremelyBuggy methods.
The end result of this age-normalization is a dataset of 691,610 methods from 92 projects, where each method has exactly 5 years of change history. All the projects in our dataset, after age normalization, as well as the number of ExtremelyBuggy methods in each project for the Balanced dataset can be viewed in our public repository. [^8] https://github.com/SQMLab/ExtremelyBuggyPublicData/blob/main/AdditionalFigures/AgeNormalizedRepositoryList.md
(a) Cumulative Distribution Function (CDF) of correlation between age and both revisions and bug-proneness for all 98 projects
(b) Percent of total methods, revisions, and bugs retained at different age-normalization thresholds.
Figure 2. Reason for selecting a 5-year threshold for age normalization. To enhance readability, we only use 11 markers when presenting each distribution; however, all data points are represented.
Both figures use the HighRecall Dataset to measure bugs
@@ 3.5. Labeling Bug Proneness
For all three datasets, we assigned each method to one of three categories based on the number of bug-fix commits in which it was involved:
@@ 3.6. Statistical Tests
After evaluating our data with the Anderson-Darling normality test (Razali et al., 2011), we found that most of the distributions in our data (e.g., the code metrics) do not follow a normal distribution. Therefore, we adopted the non-parametric Wilcoxon rank sum test to assess if there is a statistical difference between two given distributions. To understand the size of those difference, we employed the non-parametric Cliff’s d and categorized the effect sizes following Hess et al. (Hess and Kromrey, 2004): Negligible (N) ($\delta<0.147$), Small (S) ($0.147\leq\delta<0.33$), Medium (M) ($0.33\leq\delta<0.474$), and Large (L) ($\delta\geq 0.474$). Both of these statistical tests do not assume normally distributed distributions and have been widely used in software engineering research (He et al., 2015; Chen et al., 2020; Pecorelli et al., 2020; Chowdhury et al., 2019; Bangash et al., 2020). In addition, we used the non-parametric Kendall’s $\tau$ correlation coefficient instead of Pearson’s r (Inozemtseva and Holmes, 2014; Gil and Lalouche, 2017; Chowdhury et al., 2019). Unless otherwise stated, we use a significance threshold of $p\leq 0.05$ when evaluating the statistical significance of correlation coefficients.
@@ 4.1. RQ1: Proportion of ExtremelyBuggy Methods and Their Impact
Past research has shown that a high proportion of changes comes from a small segment of code (Grosfeld-Nir et al., 2007; Chowdhury, 2025). Motivated by this observation, we investigate what proportion of methods are ExtremelyBuggy and how many bugs the ExtremelyBuggy methods account for in a project.
@@
Table 1 shows the number of Buggy and ExtremelyBuggy methods for all three different datasets when all 92 projects are aggregated together. Although the high recall dataset contains a large number of ExtremelyBuggy methods (6.63%), these numbers decline significantly when we observe the HighPrecision and Balanced datasets. Specifically, only 0.04% (or 287 of the ~700000) methods are ExtremelyBuggy in the HighPrecision dataset. Another observation is that, compared to the HighRecall dataset, the ratio of Buggy to ExtremelyBuggy methods is smaller in the HighPrecision and balanced datasets. In the HighRecall dataset, approximately 37% of all methods with at least one bug-fix (Buggy and ExtremelyBuggy) are ExtremelyBuggy whereas in the HighPrecision and Balanced datasets, this percentage drops to approximately 10% and 13%, respectively.
Figure 3 presents the results at the project level. For example, in the HighRecall dataset, more than 10% of methods are buggy in 80% of the projects, and more than 10% of methods are ExtremelyBuggy in 40% of the projects. As expected, these percentages decrease substantially in the Balanced dataset and even more markedly in the HighPrecision dataset. In both datasets, less than 1% of the methods are buggy in most projects, which is even lower for the ExtremelyBuggy methods. These findings complement our earlier aggregated-level findings presented in Table 1. We also found that, for the HighPrecision dataset, ~35% of projects have no ExtremelyBuggy methods.
@@
Figure 3. CDF to show the percentages of methods for each project that are Buggy and ExtremelyBuggy. To enhance readability, we only use 11 markers when presenting each distribution; however, all 92 projects are represented.
[…]
Figure 4. CDFs of the percent of bugs covered by the ExtremelyBuggy methods for each of the three datasets for all 92 projects, excluding projects with no ExtremelyBuggy methods. To enhance readability, we only use 11 markers when presenting each distribution.
[…]
We now investigate what percent of bugs are captured by the ExtremelyBuggy methods alone. We counted the number of bugs in the ExtremelyBuggy methods in each project and divided it by the total number of bugs in the project for each of the 92 projects. To avoid overcounting the number of bugs in each project, we only count a bug once for each unique commit. The results for each dataset are presented in figure 4. For the HighRecall dataset—the most suitable dataset for this analysis, as it is likely to have fewer false negatives than the other two datasets—we observe that ExtremelyBuggy methods account for a large proportion of the bugs. For ~80% of projects, the ExtremelyBuggy methods accounted for $\geq 60\%$ of bugs, and for ~50% of projects, the ExtremelyBuggy methods covered $\geq 75\%$ of bugs. If we consider the percentage of ExtremelyBuggy methods (Table 1), this implies that, on average, only 6–7% of methods in many projects account for more than 75% of all bugs in those projects.
Not surprisingly, the results are much different for the HighPrecision and Balanced datasets. For these two datasets, we see much less bug coverage of the top percent of methods with bugs. However, this is expected as in the HighPrecision and Balanced datasets, many of the ExtremelyBuggy methods and the number of bugs they covered were excluded for maintaining higher precision.
RQ1 Summary ExtremelyBuggy methods are few in number yet account for a large share of total bugs. This is encouraging because by focusing on a small fraction of the codebase, practitioners can substantially reduce future maintenance effort. This also underscores the importance of early prediction and identification of these methods.
@@ 4.2. RQ2: Code Quality of ExtremelyBuggy Methods
Figure 5 compares the three different method classes for six of the selected code metrics using the Balanced dataset; the results are similar for all code metrics across all three datasets. Clearly, ExtremelyBuggy methods are larger, less readable, more complex, and have higher dependencies than both the Buggy and NotBuggy methods.
Figure 5. CDFs showing the distributions of different code metrics at the first commit between NotBuggy, Buggy, and ExtremelyBuggy methods for the Balanced dataset. Note that 34 of the ~700000 methods had a MaintainabilityIndex below zero. To improve readability, we excluded these methods when constructing the MaintainabilityIndex graph.
CDFs showing the distributions of different code metrics at the first commit
To confirm these visual findings, we also performed a statistical analysis to compare ExtremelyBuggy methods with both Buggy and NotBuggy methods. In Table 2, we present the results when comparing the ExtremelyBuggy methods to Buggy and NotBuggy methods grouped together for the Balanced dataset. The comparisons were done on all 92 projects’ distributions aggregated together. We can see that all the P-values are significant ($P\leq 0.05$) and that all the effect sizes are large, except for Parameters, which is medium. This indicates that there is a substantial difference between ExtremelyBuggy methods and all other methods at inception. Additionally, the sign of the effect size is negative (-) for Readability, SimpleReadability, and MaintainabilityIndex, indicating that ExtremelyBuggy methods are less readable and less maintainable. For other code metrics, the sign is positive (++), indicating that the ExtremelyBuggy methods are larger and more complex. The results are very similar for the HighPrecision dataset. However, the effect sizes for the HighRecall dataset are mostly medium. This is because the HighRecall dataset has a higher number of false positives when identifying the ExtremelyBuggy methods, resulting in more noise and less distinction of the ExtremelyBuggy methods.
Table 2. Results of the statistical significance test (Wilcoxon rank sum test) and effect size (Cliff’s delta) comparing ExtremelyBuggy methods with the combined group of Buggy and NotBuggy methods for the Balanced dataset. The method groups used in the comparison are formed by aggregating methods from all projects.
@@
In addition to the aggregated analysis, we also perform individual project analysis to obtain generalizable observations. Table 3 presents the results for the ExtremelyBuggy compared to the Buggy and NotBuggy methods grouped together for the Balanced dataset. For example, for 78.21% of projects, the difference in SLOCStandard between ExtremelyBuggy and other methods is statistically significant ($P\leq 0.05$). For these projects, the difference was large (L) for 98.36% of the projects, and for the remaining 1.64% it was medium (M). For all code metrics, some projects (usually approximately 20-30%) did not show statistical differences between ExtremelyBuggy and other methods. This is because some projects in the Balanced dataset have a very low number of ExtremelyBuggy methods, making it difficult to discern differences in distributions. However, for the projects that did have a statistical difference, that difference was almost always large. Out of the 14 code metrics, only one, Parameter, exhibited distinct characteristics. While more than 62% of projects consistently demonstrated statistical significance for all other metrics, Parameter differed for only 20% of projects, indicating that parameter count is not as influential as other metrics for identifying ExtremelyBuggy methods.
For the HighPrecision dataset, approximately 40% of projects exhibited statistical differences between ExtremelyBuggy and other methods for all metrics, with the effect size remaining almost always large. For the HighRecall dataset, approximately 90-95% of projects demonstrated statistical differences for all metrics; however, the effect size ranged from small (S) to large (L) (typically ~30% in each category). This is again due to the lower precision in the HighRecall dataset.
Table 3. The results of statistical significance test (Wilcoxon rank sum test) and effect size (Cliff’s delta) comparing ExtremelyBuggy methods with the combined group of Buggy and NotBuggy methods for the Balanced dataset, evaluated per project. Each value denotes the percentage of projects exhibiting the corresponding behavior. For instance, for the totalFanOut metric, 80.77% of projects show a statistically significant difference, and in 96.83% of those cases, the effect size is classified as large (L).
@@
To further understand how ExtremelyBuggy methods compare to Buggy and NotBuggy ones, we also did these comparisons without grouping the Buggy and NotBuggy methods together. We found that the effect size differences between NotBuggy and Buggy methods were generally small to medium. However, the difference between NotBuggy and ExtremelyBuggy methods was mostly large for all code metrics and datasets. This finding confirms the visual findings from figure 5 about the differences between the three bug-proneness types.
RQ2 Summary At their inception, ExtremelyBuggy methods are significantly larger, less readable, and more complex than other methods. This is encouraging because it suggests these methods are distinguishable at the very beginning of their lifetime.